# The consistency strength of determinacy when all sets are universally Baire

### v2

@@ 5.1. Direct limit systems at $\kappa_{0}$
Let $\delta^{*}$ and $\delta^{**}$ be cutpoints in $\mathcal{M}$ with $\delta^{*}<\delta^{**}<\kappa_{0}$. Then, in $\mathcal{W}$, $\mathcal{N}_{\delta^{*}}$ iterates to $\mathcal{N}_{\delta^{**}}$ without drops on the main branch and via the $\mathcal{P}^{0}$-to-$\mathcal{N}_{\delta^{*}}^{+}$-tail of $\Sigma_{0}$.
@@
Let $(\mathcal{L}_{\eta}\mid\eta\leq\eta^{*})$ and $(\mathcal{L}_{\eta}^{\prime}\mid\eta\leq\eta^{\prime})$ be the constructions in $\mathcal{M}$ giving rise to $\mathcal{N}_{\delta^{*}}$ and $\mathcal{N}_{\delta^{**}}$ respectively. Working in $\mathcal{W}$, compare $\mathcal{N}_{\delta^{*}}$ against the construction $(\mathcal{L}_{\eta}^{\prime}\mid\eta\leq\eta^{\prime})$, using the iteration strategy for $\mathcal{N}_{\delta^{*}}$ induced by the background strategy. By stationarity and universality of fully backgrounded constructions, see [12, Lemmas 2.11 and 2.13], $\mathcal{N}_{\delta^{*}}$ iterates to some level of the construction $(\mathcal{L}_{\eta}^{\prime}\mid\eta\leq\eta^{\prime})$ without drops on the main branch. As $\Sigma_{0}$ has branch condensation and fully backgrounded constructions absorb strategies with branch condensation (cf. [12, Lemma 2.15]) the iteration of $\mathcal{N}_{\delta^{*}}$ into a level of $(\mathcal{L}_{\eta}^{\prime}\mid\eta\leq\eta^{\prime})$ is in fact a $\Sigma_{0}$-iteration. Using that $\Sigma_{0}$ is super fullness preserving, we obtain that $\mathcal{N}_{\delta^{*}}$ iterates to $\mathcal{N}_{\delta^{**}}$. ∎
Now, we argue that we can consider $\mathcal{M}_{\infty}^{0}$ as a limit of premice of the form $\mathcal{N}_{\delta^{*}}$ for cutpoints $\delta^{*}<\kappa_{0}$ in $\mathcal{M}$. Let $\delta^{*}_{0}$ be the least cutpoint in $\mathcal{M}|\kappa_{0}$ above $\delta_{0}$. Work in $\mathcal{M}$. We will use the definitions of $\kappa_{0}$-short and $\kappa_{0}$-maximal trees $\mathcal{T}$ referring to $(Lp^{\infty}(\mathcal{M}(\mathcal{T})))^{\mathcal{W}}$ as well as other notions introduced in Section 3.2. By Lemma 5.1 these notions can be defined internally in $\mathcal{M}$. Now let $\mathcal{N}^{*}$ be the result of making $\mathcal{M}|\kappa_{0}$ generically generic over an iterate of $\mathcal{N}_{\delta^{*}_{0}}^{+}$ in the sense of [12, Discussion after Definition 3.36] via the extender algebra with many generators (see [1, Theorem 4.5]). By considering pseudo-genericity iterations in the sense of [29, Theorem 3.16] we can obtain $\mathcal{N}^{*}\in\mathcal{M}$. Recall that as $\mathcal{W}$ is a translatable structure, there is a set of term relations strongly guiding $\Sigma_{0}$. In particular, there is a set of term relations $\vec{\tau}\subseteq\mathcal{N}^{*}$ such that whenever $G\subseteq\operatorname{Col}(\omega,|\mathcal{N}^{*}|)$ is generic and $(\tau_{k}\mid k<\omega)$ is a fixed generic enumeration of $\vec{\tau}$, then $(\Sigma_{\mathcal{N}^{*}})^{G}$ is strongly guided by $(\tau_{k}\mid k<\omega)$. For $l\leq\omega$, let $f_{l}\in F(\kappa_{0},\Sigma_{0})$ be the function given by
@@
Note that the definition of $\tau_{k}^{\mathcal{Q}}$ makes sense by interpreting part of the term relation $\tau_{k}$ as a $\operatorname{Col}(\omega,((\delta^{\mathcal{Q}})^{+n})^{\mathcal{Q}})$-term relation. Then $f_{l}\in\mathcal{M}$ for every $l<\omega$. Now let
=$\mathcal{I}=\{(\mathcal{Q},l)\mid\mathcal{Q}\text{ is $\kappa_{0}$-suitable, strongly $f_{l}$-iterable and countable in }\mathcal{M}^{\operatorname{Col}(\omega,{<}\kappa_{0})}\}$
@@
Note that the models $\mathcal{N}_{\delta^{*}}^{+}$ for cutpoints $\delta^{*}<\kappa_{0}$ are cofinal in the $\Sigma_{0}$-iterates of $\mathcal{P}^{0}$ that are countable in $\mathcal{W}[g]$. By Lemmas 5.2 and 5.3 combined with standard arguments on the existence of strongly $f$-iterable premice for $f\in F(\kappa_{0},\Sigma_{0})$, $(\mathcal{I},\leq_{\mathcal{I}})$ is a directed system and, in $\mathcal{W}$, there are $\Sigma$-iterations between the models. In $\mathcal{M}$, there is, for each $(\mathcal{Q},l)\leq_{\mathcal{I}}(\mathcal{Q}^{\prime},l^{\prime})$ a natural embedding
@@
We let $\mathcal{F}$ be the directed system of models $H^{\mathcal{Q}}_{f_{l}}$ indexed by $(\mathcal{Q},f_{l})\in(\mathcal{I},\leq_{\mathcal{I}})$ together with the embeddings $\pi_{(\mathcal{Q},l),(\mathcal{Q}^{\prime},l^{\prime})}$. Let $\mathcal{M}_{\infty}^{*}$ be the direct limit of $\mathcal{F}$ and let $\pi_{(\mathcal{Q},l),\infty}\colon H^{\mathcal{Q}}_{f_{l}}\rightarrow\mathcal{M}_{\infty}^{*}$ be the corresponding direct limit embedding.
The following lemma, which is our desired internalization of $\mathcal{M}_{\infty}^{0}$, now follows as in standard HOD computation arguments, see, for example, [29, Lemma 6.32]. The second equality in Lemma 5.4 is a consequence of fullness preservation.
@@
$\mathcal{M}_{\infty}^{*}=\mathcal{M}^{0}_{\infty}=(Lp_{\omega}^{\infty})^{\mathcal{W}}(\mathcal{M}_{\infty}^{0,-})$.
Recall that $\mathcal{M}_{\infty}^{0}$ is a $\Sigma_{0}$-iterate of $\mathcal{P}^{0}$ in $\mathcal{W}$. Write $\Sigma_{\mathcal{M}_{\infty}^{0}}$ for the corresponding tail strategy of $\Sigma_{0}$ and $\pi_{\infty}^{0}\colon\mathcal{P}^{0}\rightarrow\mathcal{M}_{\infty}^{0}$ for the iteration embedding. Moreover, recall that $\mathcal{M}|\delta_{1}$ is obtained from a fully backgrounded construction in $\mathcal{W}|\delta_{1}$. The proof of the successor case in the proof of [12, Theorem 6.5] (cf. [13, Lemma 2.15]) shows that $\Sigma_{\mathcal{M}_{\infty}^{0}}$ up to $\delta_{1}$ is in $\mathcal{M}$. More precisely:
@@
$\Sigma_{\mathcal{M}_{\infty}^{0}}\upharpoonright(\mathcal{M}|\delta_{1})\in\mathcal{M}.$
We will show in Lemma 5.11 that, if we suppose inductively that $\kappa_{0}$ is fully strong in $\mathcal{M}$, then $\Sigma_{\mathcal{M}_{\infty}^{0}}\upharpoonright\mathcal{M}\in\mathcal{M}.$
@@ 5.2. Direct limit systems at $\kappa_{i}$
We now define more general direct limit systems that will be used to characterize extenders with critical point $\kappa_{i}$ for $i>0$ and argue that their direct limits $\mathcal{M}_{\infty}^{i}$ are in $\mathcal{M}$. Recall that $\kappa_{i}$ is the least ${<}\delta_{i+1}$-strong cardinal in $\mathcal{W}$ [^14] Again, by a short argument in this setting $\kappa_{i}$ will also be the first cardinal above $\delta_{i}$ that is strong up to $\delta_{i+1}$ in $\mathcal{M}$ since the fully backgrounded construction will add extenders witnessing this. As the rest of the argument in this article is structured inductively on $i<\omega$, we will assume that $\kappa_{j}$ for $j<i$ is fully strong in $\mathcal{M}$ in some of the arguments in this subsection. To avoid confusion, we will emphasize where this inductive hypothesis is used.
@@
Working in $\mathcal{W}$, let $\mathcal{M}_{\infty}^{i}$ be the direct limit of all $\Sigma_{i}$-iterates of $\mathcal{P}^{i}=Lp_{\omega}^{\Sigma_{i-1},\infty}(\mathcal{W}|\delta_{i})$ that are countable in $\mathcal{W}[g]$, where $g$ is $\operatorname{Col}(\omega,{<}\kappa_{i})$-generic over $\mathcal{W}$. By standard arguments this direct limit is well-defined and we can let $\delta_{\infty}^{i}$ be the image of $\delta_{i}$ in $\mathcal{M}_{\infty}^{i}$ under the direct limit embedding. Moreover, we write
$\mathcal{M}_{\infty}^{i,-}=\mathcal{M}_{\infty}^{i}|\delta_{\infty}^{i}.$
@@
As in the case $i=0$, we aim to identify $\mathcal{M}^{i}_{\infty}$ inside $\mathcal{M}$ as the limit of certain fully backgrounded constructions.
@@
Let $i>0$. Working in $\mathcal{M}$, we say an ordinal $\delta^{*}<\kappa_{i}$ is a $\kappa_{i}$-weak cutpoint iff all extenders $E$ on the sequence of $\mathcal{M}$ overlapping $\delta^{*}$, i.e., with $\operatorname{crit}(E)<\delta^{*}<\operatorname{lh}(E)$, have critical point $\kappa_{j}$ for some $j<i$.
@@
For any $\kappa_{i}$-weak cutpoint $\delta^{*}$ in $\mathcal{M}|\kappa_{i}$ we define $\mathcal{N}_{\delta^{*}}$ as the result of a hod pair construction [^15] Similar as in [12, Definition 2.6] but we are working in a fine structural model, not in a weak background triple. above $\delta^{*}$ in $\mathcal{M}|\kappa_{i}$ relative to certain tail strategies of $\Sigma_{k}$ for $k<i$ until it reaches $i+1$ full Woodin cardinals. We make this precise in what follows. Note that by Lemmas 5.5 and 5.11 we may and will assume inductively that the relevant tail of $\Sigma_{k-1}$ restricted to $\mathcal{M}$ is in $\mathcal{M}$ for all $k\leq i$.
@@
To keep the notation simple, we illustrate the definition of the models $\mathcal{N}_{\delta^{*}}$ at $\kappa_{1}$ and $\kappa_{2}$. The definition for $\kappa_{i}$, $i>2$, is then a straightforward generalization. In what follows, whenever we say that we “add a strategy” we mean that the strategy gets added to the model in a fine structural sense as in [12]. We decided to not go into the fine structural details here in order to be able to focus on the new ideas in this construction. We refer the interested reader to [12, Chapter 1].
@@
So let $\delta^{*}>\delta_{1}$ be a $\kappa_{1}$-weak cutpoint. Then $\mathcal{N}_{\delta^{*}}$ is given by the following construction: Perform a fully backgrounded construction as in [8] (without the smallness assumption) using extenders with critical point above $\delta^{*}$ inside $\mathcal{M}|\delta_{2}$. Let $(\mathcal{L}_{\eta}\mid\eta\leq\eta^{*})$ be the result of the construction and suppose it reaches an ordinal $\gamma$ such that
@@
If it exists, we say that $\mathcal{L}_{\gamma}$ is the level where the construction reaches the first full Woodin cardinal. We stop the construction and say that it does not succeed if no such level is reached with $\mathcal{L}_{\gamma}\cap\operatorname{Ord}<\kappa_{1}$. We will show in Lemma 5.6 that such a level $\mathcal{L}_{\gamma}$ is indeed reached below $\kappa_{1}$ and hence all extenders in $\mathcal{R}_{0}$ are backgrounded not only in $\mathcal{M}$ but also in $\mathcal{W}$.
We will argue in Lemma 5.7 that $\mathcal{R}_{0}$ is a $\Sigma_{\mathcal{M}_{\infty}^{0}}$-iterate of $\mathcal{M}_{\infty}^{0}$ (that itself is a $\Sigma_{0}$-iterate of $\mathcal{P}^{0}$). Write $\Sigma_{\mathcal{R}_{0}}$ for the respective tail strategy of $\Sigma_{0}$. Then $(\mathcal{R}_{0},\Sigma_{\mathcal{R}_{0}})$ is a hod pair. As $\Sigma_{0}$ has branch condensation, the fully backgrounded construction will absorb the tail strategy $\Sigma_{\mathcal{R}_{0}}$ by [12, Lemma 2.15]. More precisely, $\Sigma_{\mathcal{R}_{0}}$ agrees with the iteration strategy $\mathcal{R}_{0}$ inherits from $\mathcal{M}$ (and ultimately from $\mathcal{W}$) as the result of a backgrounded construction. Therefore, it makes sense to define $\mathcal{R}_{1}$ as the result of adding $\Sigma_{\mathcal{R}_{0}}$ to $\mathcal{R}_{0}$. That means, we proceed further in the fully backgrounded construction, now relative to $\Sigma_{\mathcal{R}_{0}}$, and the construction will not project across $\mathcal{R}_{0}\cap\operatorname{Ord}$. Write $(\mathcal{L}_{\eta}^{1}\mid\eta\leq\eta^{*}_{1})$ for the result of this construction. If it exists, let $\gamma_{1}$ be the least ordinal such that
@@
If it exists, we say that $\mathcal{L}^{1}_{\gamma_{1}}$ is the level where the construction reaches the second full Woodin cardinal. Again, if no such level is reached below $\kappa_{1}$ we say that the construction does not succeed. We will argue in Lemma 5.6 that the construction succeeds and in Lemma 5.8 that $\mathcal{N}_{\delta^{*}}^{+}$ is definable in $\mathcal{M}$.
@@
For each $i<\omega$ and each $\kappa_{i}$-weak cutpoint $\delta^{*}$ with $\delta_{i}<\delta^{*}<\kappa_{i}$ in $\mathcal{M}$, the construction of $\mathcal{N}_{\delta^{*}}^{+}$ does not fail, i.e., it reaches a level with $i+1$ full Woodin cardinals below $\kappa_{i}$.
@@
As discussed in the previous subsection, for $i=0$ this follows directly from [13, Lemma 2.14]. We give the argument for $i=1$ here, the general case $i>1$ is similar.
Fix a $\kappa_{1}$-weak cutpoint $\delta^{*}$ with $\delta_{1}<\delta^{*}<\kappa_{1}$ in $\mathcal{M}$ and consider the construction of $\mathcal{N}_{\delta^{*}}^{+}$. The argument that this construction reaches the first full Woodin is an easier version of the argument to follow for the second full Woodin. So suppose that the construction reaches the first full Woodin at some level $\mathcal{R}_{0}$. Then we proceed further in the fully backgrounded construction in $\mathcal{M}|\delta_{2}$, now relative to $\Sigma_{\mathcal{R}_{0}}$, and write $(\mathcal{L}_{\eta}^{1}\mid\eta\leq\eta^{*}_{1})$ for the result of this construction. We first show the following claim.

### v3

@@ 5.1. Direct limit systems at $\kappa_{0}$
Let $\delta^{*}$ and $\delta^{**}$ be cutpoints in $\mathcal{M}^{(0)}$ with $\delta^{*}<\delta^{**}<\kappa_{0}$. Then, in $\mathcal{W}$, $\mathcal{N}_{\delta^{*}}$ iterates to $\mathcal{N}_{\delta^{**}}$ without drops on the main branch and via the $\mathcal{P}^{0}$-to-$\mathcal{N}_{\delta^{*}}^{+}$-tail of $\Sigma_{0}$.
@@
Let $(\mathcal{L}_{\eta}\mid\eta\leq\eta^{*})$ and $(\mathcal{L}_{\eta}^{\prime}\mid\eta\leq\eta^{\prime})$ be the constructions in $\mathcal{M}^{(0)}$ giving rise to $\mathcal{N}_{\delta^{*}}$ and $\mathcal{N}_{\delta^{**}}$ respectively. Working in $\mathcal{W}$, compare $\mathcal{N}_{\delta^{*}}$ against the construction $(\mathcal{L}_{\eta}^{\prime}\mid\eta\leq\eta^{\prime})$, using the iteration strategy for $\mathcal{N}_{\delta^{*}}$ induced by the background strategy. By stationarity and universality of fully backgrounded constructions, see [13, Lemmas 2.11 and 2.13], $\mathcal{N}_{\delta^{*}}$ iterates to some level of the construction $(\mathcal{L}_{\eta}^{\prime}\mid\eta\leq\eta^{\prime})$ without drops on the main branch. As $\Sigma_{0}$ has branch condensation and fully backgrounded constructions absorb strategies with branch condensation (cf. [13, Lemma 2.15]) the iteration of $\mathcal{N}_{\delta^{*}}$ into a level of $(\mathcal{L}_{\eta}^{\prime}\mid\eta\leq\eta^{\prime})$ is in fact a $\Sigma_{0}$-iteration. Using that $\Sigma_{0}$ is super fullness preserving, we obtain that $\mathcal{N}_{\delta^{*}}$ iterates to $\mathcal{N}_{\delta^{**}}$. ∎
Now, we argue that we can consider $\mathcal{P}_{\infty}^{0}$ as a direct limit of premice such that densely many of these premice are of the form $\mathcal{N}_{\delta^{*}}$ for cutpoints $\delta^{*}<\kappa_{0}$ in $\mathcal{M}^{(0)}$. Let $\delta^{*}_{0}$ be the least cutpoint in $\mathcal{M}^{(0)}|\kappa_{0}$ above $\delta_{0}$. Work in $\mathcal{M}^{(0)}$. We will use the definitions of $\kappa_{0}$-short and $\kappa_{0}$-maximal trees $\mathcal{T}$ referring to $(Lp^{\infty}(\mathcal{M}(\mathcal{T})))^{\mathcal{W}}$ as well as other notions introduced in Section 3.2. By Lemma 5.1 these notions can be defined internally in $\mathcal{M}^{(0)}$. Now let $\mathcal{N}^{*}$ be the result of making $\mathcal{M}^{(0)}|\kappa_{0}$ generically generic over an iterate of $\mathcal{N}_{\delta^{*}_{0}}^{+}$ in the sense of [13, Discussion after Definition 3.36] via the extender algebra with many generators (see [1, Theorem 4.5]). By considering pseudo-genericity iterations in the sense of [31, Theorem 3.16] we can obtain $\mathcal{N}^{*}\in\mathcal{M}^{(0)}$. Recall that as $\mathcal{W}$ is a translatable structure, there is a set of term relations strongly guiding $\Sigma_{0}$. In particular, there is a set of term relations $\vec{\tau}\subseteq\mathcal{N}^{*}$ such that whenever $G\subseteq\operatorname{Col}(\omega,|\mathcal{N}^{*}|)$ is generic and $(\tau_{k}\mid k<\omega)$ is a fixed generic enumeration of $\vec{\tau}$, then $(\Sigma_{\mathcal{N}^{*}})^{G}$ is strongly guided by $(\tau_{k}\mid k<\omega)$. For $l\leq\omega$, let $f_{l}\in F(\kappa_{0},\Sigma_{0})$ be the function given by
@@
Note that the definition of $\tau_{k}^{\mathcal{Q}}$ makes sense by interpreting part of the term relation $\tau_{k}$ as a $\operatorname{Col}(\omega,((\delta^{\mathcal{Q}})^{+n})^{\mathcal{Q}})$-term relation. Then $f_{l}\in\mathcal{M}^{(0)}$ for every $l<\omega$. Now let
=$\mathcal{I}=\{(\mathcal{Q},l)\mid\mathcal{Q}\text{ is $\kappa_{0}$-suitable, strongly $f_{l}$-iterable and countable in }(\mathcal{M}^{(0)})^{\operatorname{Col}(\omega,{<}\kappa_{0})}\}$
@@
Note that the models $\mathcal{N}_{\delta^{*}}^{+}$ for cutpoints $\delta^{*}<\kappa_{0}$ are cofinal in the $\Sigma_{0}$-iterates of $\mathcal{P}^{0}$ that are countable in $\mathcal{W}[g]$. By Lemmas 5.2 and 5.3 combined with standard arguments [^20] See, for example, [31, Section 3] or [13, Section 4.1]. on the existence of strongly $f$-iterable premice for $f\in F(\kappa_{0},\Sigma_{0})$, $(\mathcal{I},\leq_{\mathcal{I}})$ is a directed system and, in $\mathcal{W}$, there are $\Sigma$-iterations between the models. In $\mathcal{M}^{(0)}$, there is, for each $(\mathcal{Q},l)\leq_{\mathcal{I}}(\mathcal{Q}^{\prime},l^{\prime})$ a natural embedding
@@
We let $\mathcal{F}$ be the directed system of models $H^{\mathcal{Q}}_{f_{l}}$ indexed by $(\mathcal{Q},f_{l})\in(\mathcal{I},\leq_{\mathcal{I}})$ together with the embeddings $\pi_{(\mathcal{Q},l),(\mathcal{Q}^{\prime},l^{\prime})}$. Let $\mathcal{P}_{\infty}^{*}$ be the direct limit of $\mathcal{F}$ and let $\pi_{(\mathcal{Q},l),\infty}\colon H^{\mathcal{Q}}_{f_{l}}\rightarrow\mathcal{P}_{\infty}^{*}$ be the corresponding direct limit embedding.
The following lemma, which is our desired internalization of $\mathcal{P}_{\infty}^{0}$, now follows as in standard HOD computation arguments, see, for example, [31, Lemma 6.32]. The second equality in Lemma 5.4 is a consequence of fullness preservation.
@@
$\mathcal{P}_{\infty}^{*}=\mathcal{P}^{0}_{\infty}=(Lp_{\omega}^{\infty})^{\mathcal{W}}(\mathcal{P}_{\infty}^{0,-})$.
Recall that $\mathcal{P}_{\infty}^{0}$ is a $\Sigma_{0}$-iterate of $\mathcal{P}^{0}$ in $\mathcal{W}$. Write $\Sigma_{\mathcal{P}_{\infty}^{0}}$ for the corresponding tail strategy of $\Sigma_{0}$ and $\pi_{\infty}^{0}\colon\mathcal{P}^{0}\rightarrow\mathcal{P}_{\infty}^{0}$ for the iteration embedding. Moreover, recall that $\mathcal{M}^{(0)}|\delta_{1}$ is obtained from a fully backgrounded construction in $\mathcal{W}|\delta_{1}$. The proof of the successor case in the proof of [13, Theorem 6.5] (cf. [14, Lemma 2.15]) shows that $\Sigma_{\mathcal{P}_{\infty}^{0}}$ up to $\delta_{1}$ is in $\mathcal{M}^{(0)}$. More precisely:
@@
$\Sigma_{\mathcal{P}_{\infty}^{0}}\upharpoonright(\mathcal{M}^{(0)}|\delta_{1})$ is amenable to $\mathcal{M}^{(0)}|\delta_{1}$, i.e., for every $X\in\mathcal{M}^{(0)}|\delta_{1}$, $\Sigma_{\mathcal{P}_{\infty}^{0}}\upharpoonright X\in\mathcal{M}^{(0)}|\delta_{1}$.
We will show in Lemma 5.12 that, if we suppose inductively that $(\mathsf{IH})_{j}$ holds, i.e., the $\mathcal{M}$-construction in $\mathcal{W}$ where Clause (2c) is restricted to extenders with critical point in $\{\kappa_{l}\mid l\leq j\}$ is well-defined up to $\delta_{j+1}$ and, if $\mathcal{M}^{(j)}|\delta_{j+1}$ is the result of this construction up to $\delta_{j+1}$, $\Sigma_{\mathcal{P}_{\infty}^{0}}\upharpoonright(\mathcal{M}^{(j)}|\delta_{j+1})$ is amenable to $\mathcal{M}^{(j)}|\delta_{j+1}$.
@@ 5.2. Direct limit systems at $\kappa_{i}$
We now define more general direct limit systems that will be used to characterize extenders with critical point $\kappa_{i}$ for $i>0$ and argue that their direct limits $\mathcal{P}_{\infty}^{i}$ are in $\mathcal{M}$. Recall that we fixed a translatable structure $\mathcal{W}$. In this subsection we assume $(\mathsf{IH})_{i}$ and let $\mathcal{M}^{(i)}$ denote the result of an $\mathcal{M}$-construction where Clause (2c) is restricted to extenders with critical point in $\{\kappa_{j}\mid j\leq i\}$. Note that $\mathcal{M}^{(i)}$ is not necessarily a proper class model at this point as the $\mathcal{M}$-construction might break down for extenders with critical point $\kappa_{i}$ added by Clause (2c) of Definition 4.3. One of the applications of the results of this subsection is that this does not happen and $\mathcal{M}^{(i)}$ is indeed a proper class model. Recall that $\kappa_{i}$ is the least ${<}\delta_{i+1}$-strong cardinal in $\mathcal{W}$ [^21] Again, by a short argument in this setting $\kappa_{i}$ will also be the first cardinal above $\delta_{i}$ that is strong up to $\delta_{i+1}$ in $\mathcal{M}^{(i)}$ since by an argument of Farmer Schlutzenberg the fully backgrounded construction will add extenders witnessing this. To avoid confusion, we will emphasize where the inductive hypothesis is used. By our inductive hypothesis, $\mathcal{M}^{(j)}$ is a well-defined proper class model for all $j<i$. Note that by construction $\mathcal{M}^{(i)}|\delta_{i+1}=\mathcal{M}^{(i-1)}|\delta_{i+1}$, so $\mathcal{M}^{(i)}|\delta_{i+1}$ is well-defined.
@@
Working in $\mathcal{W}$, let $\mathcal{P}_{\infty}^{i}$ be the direct limit of all $\Sigma_{i}$-iterates of $\mathcal{P}^{i}=Lp_{\omega}^{\Sigma_{i-1},\infty}(\mathcal{W}|\delta_{i})$ that are countable in $\mathcal{W}[g]$, where $g$ is $\operatorname{Col}(\omega,{<}\kappa_{i})$-generic over $\mathcal{W}$. By standard arguments this direct limit is well-defined and well-founded (see [13]) and we can let $\delta_{\infty}^{i}$ be the image of $\delta_{i}$ in $\mathcal{P}_{\infty}^{i}$ under the direct limit embedding. Moreover, we write
$\mathcal{P}_{\infty}^{i,-}=\mathcal{P}_{\infty}^{i}|\delta_{\infty}^{i}.$
@@
As in the case $i=0$, we aim to identify $\mathcal{P}^{i}_{\infty}$ inside $\mathcal{M}^{(i)}|\delta_{i+1}$ as the direct limit of premice where densely many of these premice result from certain backgrounded constructions, in this case relative to strategies.
@@
Let $i>0$. Working in $\mathcal{M}^{(i)}$, we say an ordinal $\delta^{*}<\kappa_{i}$ is a $\kappa_{i}$-weak cutpoint iff all extenders $E$ on the sequence of $\mathcal{M}^{(i)}$ overlapping $\delta^{*}$, i.e., with $\operatorname{crit}(E)<\delta^{*}<\operatorname{lh}(E)$, have critical point $\kappa_{j}$ for some $j<i$.
@@
For any $\kappa_{i}$-weak cutpoint $\delta^{*}$ in $\mathcal{M}^{(i)}|\kappa_{i}$ we define $\mathcal{N}_{\delta^{*}}$ as the result of a hod pair construction [^22] Similar as in [13, Definition 2.6] but we are working in a fine structural model, not in a weak background triple. above $\delta^{*}$ in $\mathcal{M}^{(i)}|\delta_{i+1}$ relative to certain tail strategies of $\Sigma_{k}$ for $k<i$ until it reaches $i+1$ full Woodin cardinals. We make this precise in what follows. This part of the argument works in $\mathcal{M}^{(i)}|\delta_{i+1}$. As already noted above, this segment of $\mathcal{M}^{(i)}$ exists by $(\mathsf{IH}.2)_{i}$ as extender with critical point $\kappa_{l}$ for $l\geq i$ that are not fully backgrounded (i.e., that are added by Clause (2c) of Definition 4.3) are only added past stage $\delta_{i+1}$ of the construction. Note that by Lemmas 5.5 and 5.12 we may and will assume inductively that the relevant tail of $\Sigma_{k-1}$ restricted to $\mathcal{M}^{(i)}|\delta_{i+1}$ is amenable to $\mathcal{M}^{(i)}|\delta_{i+1}$ for all $k\leq i$.
@@
To keep the notation simple, we illustrate the definition of the models $\mathcal{N}_{\delta^{*}}$ at $\kappa_{1}$ and $\kappa_{2}$. The definition for $\kappa_{i}$, $i>2$, is then a straightforward generalization. In what follows, whenever we say that we “add a strategy” we mean that the strategy gets added to the model in a fine structural sense as in [13]. We decided to not go into the fine structural details here in order to be able to focus on the new ideas in this construction. We refer the interested reader to [13, Chapter 1].
@@
So for the case $i=1$, let $\delta^{*}>\delta_{1}$ be a $\kappa_{1}$-weak cutpoint. Then $\mathcal{N}_{\delta^{*}}$ is given by the following construction: Until it reaches the first Woodin cardinal a hod pair construction is just a fully backgrounded construction. So we start with performing a fully backgrounded construction as in [8] (without the smallness assumption) using extenders with critical point above $\delta^{*}$ inside $\mathcal{M}^{(1)}|\delta_{2}$. Let $(\mathcal{L}_{\eta}\mid\eta\leq\eta^{*})$ be the result of the construction and suppose it reaches an ordinal $\gamma$ such that
@@
If it exists, we say that $\mathcal{L}_{\gamma}$ is the level where the construction reaches the first full Woodin cardinal. We stop the construction and say that it does not succeed if no such level is reached with $\mathcal{L}_{\gamma}\cap\operatorname{Ord}<\kappa_{1}$. We will show in Lemma 5.6 that such a level $\mathcal{L}_{\gamma}$ is indeed reached below $\kappa_{1}$ and hence all extenders in $\mathcal{R}_{0}$ are backgrounded not only in $\mathcal{M}^{(1)}$ but also in $\mathcal{W}$.
We will argue in Lemma 5.8 that $\mathcal{R}_{0}$ is a $\Sigma_{\mathcal{P}_{\infty}^{0}}$-iterate of $\mathcal{P}_{\infty}^{0}$ (that itself is a $\Sigma_{0}$-iterate of $\mathcal{P}^{0}$). Write $\Sigma_{\mathcal{R}_{0}}$ for the respective tail strategy of $\Sigma_{0}$. Then $(\mathcal{R}_{0},\Sigma_{\mathcal{R}_{0}})$ is a hod pair. As $\Sigma_{0}$ has branch condensation, the fully backgrounded construction will absorb the tail strategy $\Sigma_{\mathcal{R}_{0}}$ by [13, Lemma 2.15]. More precisely, $\Sigma_{\mathcal{R}_{0}}$ agrees with the iteration strategy $\mathcal{R}_{0}$ inherits from $\mathcal{M}^{(1)}$ (and ultimately from $\mathcal{W}$) as the result of a backgrounded construction. Therefore, it makes sense to define $\mathcal{R}_{1}$ as the result of adding $\Sigma_{\mathcal{R}_{0}}$ to $\mathcal{R}_{0}$. That means, we proceed further in the fully backgrounded construction, now relative to $\Sigma_{\mathcal{R}_{0}}$, and the construction will not project across $\mathcal{R}_{0}\cap\operatorname{Ord}$. Write $(\mathcal{L}_{\eta}^{1}\mid\eta\leq\eta^{*}_{1})$ for the result of this construction. If it exists, let $\gamma_{1}$ be the least ordinal such that
@@
If it exists, we say that $\mathcal{L}^{1}_{\gamma_{1}}$ is the level where the construction reaches the second full Woodin cardinal. Again, if no such level is reached below $\kappa_{1}$ we say that the construction does not succeed. We will argue in Lemma 5.6 that the construction succeeds and in Lemma 5.9 that $\mathcal{N}_{\delta^{*}}^{+}$ is definable in $\mathcal{M}^{(1)}|\delta_{2}$.
@@
For each $i<\omega$ and each $\kappa_{i}$-weak cutpoint $\delta^{*}$ with $\delta_{i}<\delta^{*}<\kappa_{i}$ in $\mathcal{M}^{(1)}$, the construction of $\mathcal{N}_{\delta^{*}}^{+}$ does not fail, i.e., it reaches a level with $i+1$ full Woodin cardinals below $\kappa_{i}$.
@@
As discussed in the previous subsection, for $i=0$ this follows directly from [14, Lemma 2.14]. We give the argument for $i=1$ here, the general case $i>1$ is similar.
Fix a $\kappa_{1}$-weak cutpoint $\delta^{*}$ with $\delta_{1}<\delta^{*}<\kappa_{1}$ in $\mathcal{M}^{(1)}$ and consider the construction of $\mathcal{N}_{\delta^{*}}^{+}$. The argument that this construction reaches the first full Woodin is an easier version of the argument to follow for the second full Woodin. So suppose that the construction reaches the first full Woodin at some level $\mathcal{R}_{0}$. Then we proceed further in the fully backgrounded construction in $\mathcal{M}^{(1)}|\delta_{2}$, now relative to $\Sigma_{\mathcal{R}_{0}}$, and write $(\mathcal{L}_{\eta}^{1}\mid\eta\leq\eta^{*}_{1})$ for the result of this construction. We first show the following claim.